Can you help me please to solve this equation:The sum of three consecutive integers is 78.Find the number.
Thanks

Answer by jayanthihemkos@yahoo.com(87)

The sum of three consecutive integers is 78.Find the number.
Let the first integer be = x
second integer = x+1
third integer = x+2

sum of the integer
x+(x+1) +(x+2) = 78
3x+3 = 78
3x +3-3 = 78-3
3x = 75
3x/3 = 75/3
x = 25
x+1 = 25+1 = 26
x+2 = 25 +2 = 27
so the integers are 25,26,27
check
25+26+27 = 78